OpenStudy (anonymous):

In the function y=-3x^2+1 , what effect does the negative sign have on the graph, as compared to the graph of y=x^2? A. It reflects the graph across the x-axis. B. It shrinks the graph vertically by a factor of 3. C. It stretches the graph horizontally by a factor of 3. D. It shrinks the graph horizontally by a factor of 3. ***Idk what it means though.. :/ please explain? thx :)

OpenStudy (anonymous):

Think about if you were to plot a value that will be far away from the x-axis, like 3: y=3(3)^2+1 = 10 but when you put that negative on: y=-3(3)^2+1 = -8 It will flip (or reflect) every result over the x-axis

OpenStudy (anonymous):

ohhh so my answer is A. it reflects the graph across the x axis?

OpenStudy (anonymous):

You betcha

OpenStudy (anonymous):

haha thanks :) and i understand the explanation too :) thanks @tonks :) and i like ur username :) its like tonks from harry potter!! haha :P

OpenStudy (anonymous):

It's also my name, for real

OpenStudy (anonymous):

haha cool!! :P